Instrument Log
TSI-880(660)            C1       Central Facility, Lamont, OK

Date

Technician

Comments

6/9/2016

MS

Notice that the strip was not blocking the sun during PM rounds this morning.  Tried to restarted the program twice but to no avail,  then went to the instrument and shut the power off at 1512 GMT, and the breaker off.  Then disconnected both wire from the instruments battery.  Reversed actions and opened up mirtimeHT to let it take off, then opened up HSI retriever to bring instrument back up.  Strip is shading now.   RTS at 1527 GMT.

5/31/2016

MS

Strip was not shading the mirror and also the strip was facing due west.  Stopped program and opened MirtimeHT.  After the system ran through its systems check, I opened the HSI Retriever up and system is running and shading correctly now.  RTS at 1542 GMT.

4/22/16

Mentor

Mentor polished mirror from 1647 to 1656 GMT.

4/22/2016

MS

Noticed that there had been no images being taken from the TSI.  Disconnected the camera cable and then reconnected it at 1415 GMT.  RTS at 1430.

3/21/2016

MS

Strip was not shading the sun from the mirror and the program was not running due to a power surge this morning at 1235 GMT.  At 1439 GMT closed out and restarted program.  RTS 1454 GMT.

2/16/2016

MS

Program was not running at 1455 GMT, 2/16/2016.  Restarted instrument program at 1455 GMT,  RTS at 1510 GMT, 2/16/2016.

1/27/2016

MS

Instrument is down as of 1603 GMT, 1/26/2016 for a scheduled Virtual Computer Host upgrade.  Data will be lost until upgrade is complete.

12/31/15

JEM

The Central Facility lost power due to an ice storm.  The TSI did not resume normal operation after power was restored.  Troubleshooting found the power supply and voltage regulator have failed.  R/R the power supply and voltage regulator.  Removed ice from TSI.  Connected to the TSI and reset the date, time, latitude and longitude.  Checked alignment and it was observed OK.  RTS @1500 GMT.

12/15/15

MS

Computer needed to be restarted following scheduled software and security patch updates.  At 1404 GMT computers were updated and rebooted by the IT department, some data might be lost.

11/17/15

MS

A 7 hour power outage starting at 0700 GMT and power was returned at 1435 GMT, 11/17/2015. Instrument was down. During that time some data might be lost?  Started program back up and RTS at 1525 GMT, 11/17/2015.

6/10/2015

MS

Shutdown TSI program at 1347 GMT due to a scheduled power outage.  Restarted program at 1539 GMT.  RTS 1554 GMT.

6/4/2015

MS

Shutdown TSI program at 1316 GMT due to a scheduled power outage.  Restarted program at 1839 GMT.  RTS 1854 GMT.

12/15/14

CAM

The AC circuit breaker for TSI power was inadvertently left off following maintenance on Friday, 12/12.  Contacted site operator on morning of 12/13 and had the breaker turned back on.  Monday morning status check found the instrument batteries had apparently died before the breaker was turned back on as the program had stopped capturing images at 2248 GMT on 12/12.  Restarted program and RTS at 1420 GMT on 12/15.

12/12/14

CAM/JEM

The MWR was recently relocated and therefore its old stand needed to be removed.  Power and data cables for the TSI originally ran through a Hoffman enclosure on the stand so they needed to be rerouted before the stand could be removed.  Rerouted the TSI power and data cables between 1550 and 2100 GMT.  TSI data may be missing and/or corrupt during this time period.  Note that the media converter was also relocated from the old MWR enclosure to the TSI instrument enclosure.

5/14/14

CAM

No data had been collected from TSI since 1902 GMT on 5/13.  Logged in remotely and found several java exception errors.  The first was a denial of access at C:\Program Files (x86)\HSI Retriever\hsi.properties.  There was also a taskeng.exe window open indicating “Error with config file armtsimvbat.cnf.  This file must be in the same directory as “armtsivmbat.exe”.  Noted that 20140513 data directory was empty, but 20140514 directory had data up to 0124 GMT.  Suspect errors are related to Windows 7 upgrade that occurred yesterday.  Will notify IT department.  Cleared all errors and restarted HSI Retriever at 1208 GMT.  Image capture resumed status O.K.

5/13/14

TMG

Virtual machine OS upgraded to Windows 7.

5/6/14

V. Morris

Cleaned mirror at 1754 GMT.  Changed sky filter thresholds from c/t=55 and t/0=70 to 50 and 75, respectively. 

5/6/14

JEM

The TSI camera alignment is off after yesterday’s maintenance.  The western edge of the mirror is clipped off.  Adjusted the mirror alignment between 1300 GMT and 1500 GMT.  The mirror is now centered in the camera view.  RTS @1500 GMT.

5/5/14

V. Morris

Replaced TSI mirror at 1909 GMT while instrument was down for fan and control board repair.  Restarted TSI and RTS at 2142 GMT.

5/5/14

JEM

The external TSI fan is not working and the mentor would like to add an internal fan as well.  Troubleshooting found that the 12V wire had broken behind the external fan.  Repaired this wire and reinstalled the fan.  Also mounted another 12V fan on the inside of the closure blowing across the heat sink fins and connected it directly to the 12V battery so that it will run all the time circulating air inside the enclosure.  During this maintenance communication with the camera was lost.  Troubleshooting found that the 2A soldered in fuse on the control board had failed.  Shut down the TSI and removed the control board for repair in the repair lab.  Replace the 2A camera fuse and also discovered that the 2A heater relay fuse was bad so it was replaced as well.  After reinstalling the the board and powering up the TSI communication with the camera could still not be established.  Further troubleshooting found a bad connection on the 12V plug that connected to the control board.  Fixed this connection and verified 12V at the camera.  Communication could still not be established and further troubleshooting found a bad connection on the camera ethernet cable.  After this was corrected, communication with the camera was restored.  The mentor also replaced the TSI mirror during this maintenance.  TSI image collection was restarted at 2142 GMT.  RTS @2200 GMT.

5/3/14

V. Morris

Shut TSI down at 1854 GMT and replaced chrome mirror.  Restarted image capture software at 1927 GMT.  Centered camera over mirror at 2038 GMT.  Cleaned mirror and camera window and RTS at 2042 GMT.

4/10/14

JEM

The TSI has lost its settings several times recently after power bumps or power outages.  Checked to see if the systems AC outlet was on UPS power and it was not.  Checked the voltage of the internal battery and it was 0.5V.  Checked the regulator battery output voltage and it was 0V.  A close look showed that the charge regulator had been bypassed and the TSI was running on straight AC power only which would cause the TSI to stop every time there was a power bump.  Verified the regulator was operational and it was.  Reconnected the 15V power supply to the "solar" terminals on the regulator.  Reconnected the 12V battery to the battery terminal of the regulator.  Reconnected the TSI to the load terminals of the regulator.  Also reworked the AC power cable for the heater to remove unnecessary splices.  RTS @1600 GMT.

3/12/14

PAD/KT/JEM

Shadowband is not aligned.  Remotely accessed computer.  Opened TSI_Lan but no values are shown.  Ran Mirtime _HT program but it did not correct the alignment issue.  It appears Mirtime program is communicating with TSI control board.  Re-opened TSI_Lan but still no values are displayed.  Powercycled instrument.  TSI_Lan program now shows values but date had reverted to default.  Ran Mirtime program and shadowband alignment is now correct.  RTS at 1820 GMT.

3/10/14

KT

The alignment of the sunstrip needed to be adjusted. The clock on the mirror controller was 1hr. and 15min. off and needed to be adjusted. After adjustments to the clock were made the sunstrip rotated and began blocking the sun. RTS at 1345 GMT.

2/3/14

JEM

DQPR #3983 requests that the camera be adjusted to be better centered on the mirror.  Adjusted the camera housing and camera between 1600 GMT and 1640 GMT.  RTS @1645 GMT.

10/11/13

JEM

The TSI was reinstalled yesterday, but set up needs to be completed.  There was a broken ground wire on the 25 pin connector for the encoder.  Repaired this connection.  The 25 pin connector for the lantronix device had been changed to a 9 pin connector.  Used a DB25-DB9 adapter and DB9-DB9 gender changer to connect the (in case the 9 pin is ever needed again this will make it easier the changing out connectors) lantronix device.  Set up the lantronix device with the proper IP address and port configuration.  Set up the camera with the proper IP address.  Set the time and location information and then made a slight adjustment to mirror alignment.  Started the collection process and notified the mentor that the instrument was operational @1850 GMT.

10/10/13

CAM

The TSI that was loaned out for MAGIC campaign has returned and needs to be swapped out with the AMF2 TSI that was temporarily installed on 6/7/12.  Stopped program and powered instrument off at 1830 GMT.  R/R *TSI.  Replacement TSI had been modified for GPS use on the ship during MAGIC.  Returned TSI to its normal configuration for operation at SGP.  Mentor is out of the office, but will return tomorrow and needs to change network and software settings before instrument is ready for operation.  TSI remains offline at this time.

*Old TSI 660 - SN 104, WD 59108.

*New TSI 660 - SN 106, WD 30881. 

2/26/13

CAM

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Powered optical cluster UPS off at 1730 GMT.  Power was restored at 1920 GMT.  Restored UPS power at 1945 GMT. Instrument RTS status O.K and data collections resumed once SDS operations were restored.

2/4/13

PAD

Remotely accessed computer at 1412 GMT.  HSI_Retriever program was not running.  Started program.  RTS at 1427 GMT.

1/28/13

PAD

HSI_Retriever program is not running.  Started program.  RTS at 1414 GMT.

1/22/13

PAD

Computer is on but HSI_Receiver program is not running.  Started program.  RTS at 1452 GMT.

12/31/12

CAM

Unable to RAdmin into TSI virtual computer.  Windows administrator logged into virtual computer and found that HSI Retriever program had encountered a problem.  Virtual computer was trying to reboot but error message was preventing it.  Closed error message and computer rebooted at 1600 GMT.  RTS status O.K. at 1602 GMT.

12/11/12

CAM

Last successful SDS collection was at 1802 GMT on 12/10.  Attempted to RAdmin into TSI computer but was unable to.  Computer would respond for login credentials, but would not bring up the desktop once logged in.  Sent remote restart command at 1346 GMT.  Was able to successfully log in after reboot.  Restarted TSI software and started image capture.  RTS at 1351 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  TSI was gracefully shut down at 1347 GMT in anticipation of power outage.  Power was restored at 2115 GMT after electrical upgrades were complete.  Restarted TSI and RTS at 2125 GMT.

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  TSI computer displayed error: jave.net.ConnectException: connection timed out.  Cleared error and restarted program at 2130 GMT.

10/16/12

CAM

DQPR 3540 states the enclosure of the new camera allows rainwater to collect on the surface of the recessed glass window, distorting the image. To prevent this, please form a drip edge around the bottom of the enclosure with thick electrical or splicing tape.  Formed a drip edge around the bottom of the enclosure using tape as suggested.  RTS at 1705 GMT.

6/7/12

JEM

Mentor is on site this week performing upgrades to instrument.  R/R Yankee Environmental Services *TSI.  Unable to communicate with instrument due to wrong software loaded on Lantronix.  On 6/11/12, at 2020 GMT, directly connected with Lantronix serial port and was able to enter IP address.  Instrument is now communicating with computer.  RTS at 2045 GMT. 

*Old TSI 660 - SN 106, WD 30881. 

*New TSI 660 - SN 104, WD 59108.

5/23/12

PAD

At 1253 GMT, observed Image Capture Error shown on TSI program.  Troubleshooting found UPS located at base of MWRC1 has an A4 error and needs to be replaced.  Stopped program and R/R Best 660 UPS.  Restarted program.  RTS at 1533 GMT.

5/21/12

KT/PAD

Mirror shadowband is shading but needs adjustment.  Adjusted shadowband at 1448 GMT.  During adjustment computer lost comms with instrument.  Cause of outage is unknown.  Comms were reestablished within 5 minutes and shadowband adjustment was continued.  Found mirror was not tight on base.  Tightened mirror and finished shadowband adjustment.  RTS at 1518 GMT.

5/15/12

PAD

Shadowband is shading but needs adjustment.  Adjusted shadowband +2 degrees at 1445 GMT.  RTS at 1500 GMT.

5/7/12

PAD

Image Capture Error shown:  java. net. Connect Exception.  Connection Timed Out.  Stopped/ restarted program with same error shown.  Unable to connect with Hyperterminal.  Troubleshooting with network administrator found ethernet/ fiber media converter had failed.  R/R media converters with newer versions.  Comms were reestablished.  RTS at 1533 GMT.

5/1/12

PAD

Storm moved through CF area last night.  Shadowband is not shading correctly.  At 1350 GMT, adjusted shadowband 20 degrees.  RTS at 1405 GMT.

4/16/12

PAD

TSI shadowband needs aligned.  At 1329 GMT, moved shadowband +4 degrees and reset zero.  Shadowband is now in correct alignment with sun.  RTS at 1344 GMT.

2/27/12

PAD

TSI shadowband is shading but needs adjustment.  Adjusted shadowband +3 degrees at 1618 GMT.  RTS at 1633 GMT.

12/15/11

PAD

Shadowband is not completely shading sun.  Adjusted shadowband +3 degrees.  RTS at 1610 GMT.

8/24/11

PAD

Shadowband is not shading.  Adjusted shadowband -3 deg at 1334 GMT.  After cleaning mirror at 1438 GMT, shadowband was again out of alignment.  At 1442 GMT, tightened mirror on base and readjusted shadowband -36 deg.  After checking alignment, rotated mirror a few times.  Shadowband is still shading correctly.  RTS at 1457 GMT.

8/23/11

PAD

Shadowband is not shading correctly.  Adjusted shadowband +5 deg at 1729 GMT.  RTS at 1744 GMT.

8/9/11

PAD

CF lost power on 8/8/11 at 2247 GMT.  Power was restored on 8/9/11 at 0211 GMT.  At 1310 GMT, TSI Manager program was running, shadowband was not shading.  Adjusted shadowband +4 deg.  Instrument is producing data but SDS is not finding it.  Notified mentor.  Suspect certain scripts need to run that would normally run during time computer was down.  RTS at 1325 GMT.

8/1/11

PAD

While cleaning mirror at 1416 GMT, found mirror was not tight on base.  Tightened mirror on base at 1416 GMT.  Remotely accessed TSI computer at 1533 to check alignment.  Adjusted shadowband -5 deg.  RTS at 1548 GMT.

7/13/11

PAD

ADAM 4 port, model 6510 ethernet hub located in TSI enclosure was damaged during storms that came through CF on 7/3/11.  Temporary hub was installed on 7/7/11.  New replacement hub needs to be installed.  Shut off power at 1922 GMT and installed replacement hub.  Powered up system.  At 2120 GMT, RA virtual computer and found shadowband was not shading correctly.  Found control board had lost correct time, position, etc.  Ran Mirtime HT program and shadowband moved into correct position.  RTS at 2125 GMT.

7/7/11

PAD/CAM

Storms had moved through the area and SDS stopped collecting data sometime around 2300 on 7/3/11.  Network hub was damaged in Optical trailer but was still unable to communicate with camera or circuit board after hub was bypassed.  Trouble shooting found damaged ethernet hub located in TSI enclosure.  Network administrator installed temporary ethernet hub.  Now able to communicate with camera and circuit board.  RTS at 1400 GMT.

6/29/11

PAD

Adjusted shadowband -2 degrees at 1405 GMT.

6/21/11

PAD

Adjusted shadowband +2 deg at 1312 GMT.  RTS at 1327 GMT.

6/16/11

PAD

Shadowband is not shading correctly.  Adjusted shadowband +13 deg at 1341 GMT.  RTS at 1356 GMT.

5/24/11

PAD

Shadowband needs adjustment.  Adjusted shadowband +3 deg.  RTS at 1819 GMT.

3/15/11

PAD

Shadowband is shading but needs adjustment.  Adjusted shadowband 2 deg at 1636 GMT.  RTS at 1651 GMT.

3/4/11

JEM

The computer has powered off overnight again.  The computers power light is flashing amber which generally indicates power is being received, but an internal power problem exists.  Due to the multiple capacitors that are swelled and leaking on the motherboard, complete system failure is inevitable.  Powered the system off @1530 GMT and took the computer to the computer lab.  Tim Grove is working on building a virtual machine to run the TSI.  Instrument will be down until completion of the virtual machine.  Did not RTS at this time.

3/3/11

PAD

Computer and monitor power indicator lights are flashing.  Computer power supply was replaced on 3/2/11 but that did not correct the problem.  Assisted computer department and opened computer at 1412.  Observed several capacitors on computer board swelled indicating heat or possible electrical damage.  Computer department is consulting with mentor concerning requirements for replacement computer.  Program started and computer is operating at this time.  RTS at 1427 GMT.

3/2/11

PAD

Computer power supply cooling fan has failed.  Shut down computer and took to computer department.  R/R power supply.  Installed computer, powered up and started TSI Manager program.  RTS at 1504 GMT.  Computer model DHP, SN DN34K51.

3/1/11

PAD

Computer and monitor power indicator lights are flashing.  Power cycled computer at 1433 GMT.  Computer and TSI Manager program started correctly.  RTS at 1448 GMT.

2/28/11

PAD

Computer and monitor power indicator lights are flashing.  Power cycled computer at 1536 GMT.  Computer and TSI Manager program started correctly.  RTS at 1551 GMT.

2/22/11

PAD

TSI Manager shows no raw or processed images.  Bottom task bar shows Night and 264 images that have not transferred.  Used Task Manager to close out TSI Manager program.  Restarted TSI Manager and stored images began to transfer after restart.  RTS at 1607 GMT.

2/7/11

PAD

Last SDS collection was on 2/6/11 at 0203.  On/off switch on computer is blinking yellow.  Powercycled computer and it restarted correctly.  Started image capture on TSI Manager program.  Computer will communicate with camera and images are now being taken.  Shadowband is in correct alignment.  RTS at 1554 GMT.

12/1/10

CAM

Site power needs to be shut off to install the transformer and run power to the new office building.  Instruments need to be shut down gracefully.  Stopped the program and powered down at 1428 GMT.  Restarted after power was restored and RTS at 1829 GMT.

11/18/10

CAM

AC power needed to be secured to drill pier holes for new construction near the RCF trailer.  Need to gracefully shut the instrument down prior to power outage.  Powered instrument down at 1433 GMT.  Restarted instrument after AC power was restored.  RTS at 1725 GMT.

11/10/10

CAM

AC power will be interrupted for construction near the main compound.  Need to gracefully shut the instrument down prior to power outage.  Powered instrument down at 2147 GMT.  Restarted instrument after AC power was restored.  RTS at 2322 GMT.

10/7/10

PAD

Adjusted shadowband +2 degrees at 1452 GMT.

9/20/10

PAD

Last SDS collection was at 2102 GMT.  Raw and processed images are not shown on TSI Manager program.  Status bar shows "Night" and 499 images that did not transfer.  Used Task Manager to stop TSI Manager program then restarted.  After restart, images began to transfer and current mirror position is now correct.  RTS at 1433 GMT.

8/3/10

PAD/CAM

System has been down since 8/2/10 at 1808 GMT.  Previous troubleshooting has not found the problem described on CM from that date.  Further troubleshooting indicates problem with external fan and camera not getting power is still a control board issue.  Installed CPU from board #205 on board #206 with no change.  Removed board #206 and installed board #205 with its CPU.  Instrument was active when powered up.  Now able to communicate with board and camera.  Ran Mirtime program and instrument is now shading correctly.  RTS at 1841 GMT.

8/2/10

PAD/CAM

When connected to instrument with Hyperterminal, TT8 temperature is showing temperatures up to 71C during the day.  Mentor requested fan placed in instrument enclosure to move air across board to see if temperature can be lowered.  Connected 12V fan to battery located in enclosure.  Immediately after, observed external fan stopped.  I was then unable to connect with TT8 board or camera.  Board fuses are good but no power is getting to camera or external fan.  Troubleshooting indicated a problem with the board.  Consulted with mentor.  Removed board #203 and installed board #206 with CPU from board #203.  No change observed.  Troubleshooting at this time.  RTS at 2215 GMT.

7/30/10

PAD/JM

Enclosure heat sink fan is not operating.  Troubleshooting found blown fuse on control board that sends power to fan.  James Martin and I removed board and took to lab where fuse was replaced.  After powering up instrument, fan started.  Warning is still being given when using Hyperterminal to connect with instrument:  Warning.  TT8 Temperature High! **.  Reset zero and verified instrument is shading correctly.  RTS at 2130 GMT.

7/26/10

PAD

Last SDS collection was at 0102 GMT.  No raw or processed images are shown on TSI Manager program.  Task bar on bottom on TSI Manager show night and 205 images that have not transferred.  Remotely accessed computer and tried to close out/restart TSI Manager.  Program was unresponsive.  Rebooted computer and program started correctly and transferred images after restart.  RTS at 1340 GMT.

7/15/10

PAD

TSI enclosure cooling fan is not working.  Consulted mentor who requests voltage checked at board that supplies power to fan.  No voltage at J5 connector on board.  AC/DC converter located inside enclosure has 120V AC on input side and .28V DC on output side.  Wires connect converter to J5 connector on board.  With J5 connector unplugged, 0V DC is shown on converter output.  Notified instrument mentor.  Waiting on instructions.  RTS at 1447 GMT.

6/11/10

PAD

Last SDS colletion was at 1403 GMT.  There are 343 images that have not transferred from TSI Manager.  Stopped/ restarted TSI Manager.  Images started to transfer after restart.  RTS at 1719 GMT.

5/17/10

PAD

CF lost power on 5/16/10 between 0952-1504 GMT.  Powered up computer at 1402 GMT.  Image Capture error shown after startup.  Reset Axis IP camera SN and IP address.  Images are now being collected but shadowband is not aligned.  Ran mirtime.exe program to reset camera.  Shadowband is now shading.   SDS still is not collecting data from computer.  Notified mentor.  T/S found scheduled tasks were not performed while computer was down.  Mentor forced scheduled processing tasks to run and moved files from processed folder to ingest folder.  Data is now being collected by SDS.  RTS at 1417 GMT.

5/13/10

PAD

CF lost power from 0623-0759 GMT.  At 1348 GMT, image capture error is shown on monitor.  Tried to reset Axis IP according to instructions.  Message shown:  IP address is already in use.  Image is now shown but mirror is constantly rotating.  Connected to instrument with Hyperterminal, several values listed are out of range.  Notified instrument mentor who R/A into computer and ran mirtime.HT.exe program.  Correct information was sent to instrument and it is now operating normally.   RTS at 1545 GMT.

4/20/10

PAD

Adjusted shadowband -1 degrees at 1432 GMT.  RTS at 1447 GMT.

4/16/10

V. Morris

Instrument mentor on site and replaced unmodified Rev. B control board (SN 206) with new board (SN 203).  TT8 #1007221 was transferred from the old board to the new and its firmware upgraded and tested.  Boards SN 205 and SN 206  were modified (RTC circuit removed) and tested but 205 did not pass the temperature test.  

4/7/10

PAD

Shadowband is not in correct alignment.  Tightened mirror on base.  Aligned shadowband and reset 0.  RTS at 1517 GMT.

4/6/10

PAD

Adjusted shadowband - 2 degrees at 1444 GMT.  RTS at 1459 GMT.

3/29/10

PAD

Adjusted shadowband +3 degrees at 1435 GMT.  RTS at 1450 GMT.

3/24/10

CAM

Opened lid at 1943 GMT to record interface board SN.  RTS at 1950 GMT.

3/9/10

PAD

Adjusted shadowband -3 degrees at 1529 GMT.

3/2/10

PAD

Shadowband is shading but needs adjustment.  Adjusted shadowband -3 degrees.  RTS at 1615 GMT.

2/23/10

PAD

Shadowband needs adjustment.  Adjusted shadowband -3 degrees at 1648 GMT.  RTS at 1703 GMT.

2/11/10

PAD

Last SDS collection was at 2002 GMT.  There are 296 images on TSI manager that have not transferred.  Stopped/ restarted TSI Manager.  Images began to transfer.  RTS at 2311 GMT.

2/8/10

PAD

Last SDS collection was at 1502 GMT.  Remotely accessed computer and TSI manager shows 90 images that have not transferred.  Stopped/ restarted TSI manager.  Images transferred.  RTS at 1628 GMT.

2/1/10

PAD

TSI Manager shows 81 images that have not transferred.    Stopped/ restarted TSI Manager.  Images started transferring after restart.  RTS at 1554 GMT.

1/19/10

PAD

Last SDS collection was at 0102 GMT.  Status bar on TSI Manager shows night and 373 images that have not transferred. Stopped/ restarted TSI manager program.  Images started transferring.  RTS at 1458 GMT.

1/6/10

PAD

TSI manager task bar shows 278 images.  Stopped/ restarted TSI Manager.  Images started transferring after restarting program.  RTS at 1702 GMT.

12/28/09

PAD

Shadowband is only partially shading sun.  Recent rain and snow may have restricted movement of mirror.  Adjusted mirror +7 deg at 1518 GMT.  RTS at 1533 GMT.

10/19/09

PAD

Shadowband is shading but needs adjustment.  Adjusted shadowband +2 degrees at 1434 GMT.  RTS at 1449 GMT.

9/16/09

PAD

Lasts SDS collection was on 8/15/09 at 1803 GMT.  Able to remotely access computer and program is running correctly.  Rebooted computer and started TSI manager program after reboot.  SDS successfully collected at 1303 GMT.  RTS at 1303 GMT.

8/24/09

PAD

Shadowband is blocking sun but needs adjustment.  At 1448 GMT adjusted shadowband +2 degrees and reset 0.  Shadowband is now in correct alignment.  RTS at 1503 GMT.

8/4/09

V. Morris

Remotely optimized times of scheduled tasks at 1700 GMT.

7/27/09

PAD

SDS stopped collecting at 1902 GMT.  At 2052 remotely accessed computer and it shows 164 images.  Stopped and restarted program.   Images transferred and SDS successfully collected at 2102 GMT.  RTS at 2107 GMT.

7/22/09

PAD

Shadowband is still shading but needs adjustment.  Adjusted +3 degrees at 1417 GMT.  RTS at 1432 GMT. 

7/2/09

PAD

Adjusted shadowband +2 degrees at 1452 GMT. 

7/1/09

PAD

UPS that supplies power to TSI control board is shut down.  R/R UPS located at MWR base at 1224 GMT.  Started TSI Manager program.  RTS at 1239 GMT.

6/29/09

PAD

Last SDS collection was at 0102 GMT.  Found UPS that supplies power to TSI control board was off.  Power cycled UPS and restarted TSI Manager program.  Tested UPS with no problems found.  RTS at 1442 GMT.

6/25/09

TMG

Computer software needs updated.  Stopped image capture @ 1638 GMT.  Installed Service Pack 3 @ 1639 GMT.   Rebooted and RTS @ 1722 GMT.

6/23/09

PAD

Last SDS collection was at 0403 GMT.  Remotely access computer and error was shown: Java.net. NoRouteToHostException.  Operation timed out.  Program failed to restart.  Found UPS that supplies power to instrument control board shut down.  Power cycled UPS and started TSI Manger program.  RTS at 1253 GMT.

6/11/09

PAD

Last SDS collection was at 0402 GMT.  Neither raw nor processed images are shown on TSI Manager.  Status shows:  Waiting.  55 images.  Stopped/restarted TSI Manager.  After restart over 1000 images were shown.  After stored images were transferred, monitor shows current image.  RTS at 1408 GMT.

6/10/09

PAD

Enclosure cooling fan is not working.  Troubleshooting has determined problem with instrument control board.  R/R instrument control board at 1511 GMT.  Restarted program and checked mirror shading alignment.  Cooling fan is now working correctly.  Old SN 202.  New SN 206.  RTS at 1541 GMT.

6/8/09

PAD

Stopped program at 1434 GMT.  Polished mirror and restarted program at 1448 GMT. 

5/29/09

PAD

Instrument enclosure cooling fan is not working.  Program shows cooling fan is on but it is not running.  Checked voltage at fan and found no voltage.  Contacted mentor to troubleshoot issue.  Mentor stated power for fan comes from control board located in enclosure and indicates it could be blown fuse on board.  Troubleshooting at 1950 GMT found no voltage at board going to fan. Troubleshooting at this time.  RTS at 2015 GMT.

5/29/09

PAD

Status bar shows 85 images.  Data is collected but not transferred when this happens.  Raw and processed images are shown but are not current.  Stopped Task Manager, exited, and restarted.  Images started transferring but stopped again at 20.  Stopped program and rebooted computer.  Started program and all remaining images were transferred.  RTS at 1543 GMT.

5/19/09

PAD

Last successful collection was on 5/19/09 at 1902 GMT.  TSI manager is on monitor, but raw and processed images shown are not current.  Status bar shows images being taken every 30 seconds and 190 total images.  Stopped and restarted TSI manager.  Current image is now displayed.  RTS at 2110 GMT.

5/18/09

PAD

Last successful collection was on 5/17/09 at 1902 GMT.  TSI manager is on monitor, but raw and processed images are not shown.  Printed screen and saved to desktop.  Status bar shows:  Night  878.  TSI manager will not close out.  Used Task Manager to end program.  Restarted program and 878 images started transferring.  Current image is now displayed.  RTS at 1423 GMT.

5/13/09

PAD

TSI shadowband is still shading but needs to be adjusted.  Adjusted shadowband 4 degrees at 1535 GMT.  RTS at 1550 GMT.

4/1/09

PAD

At 1417 GMT, display status shows night and 323 images.  No raw or processed images are shown.  TSI Manager would not close out.  I had to use task manager to close program.  Rebooted computer and restarted TSI Manager.  Program started correctly.  RTS at 1432 GMT.

10/17/08

PAD

Shadowband is still shading mirror but needs adjustment.  Adjusted shadowband 3 degrees at 1513 GMT.  RTS at 1528 GMT.

9/10/08

PAD

Observed at 1431 GMT, raw image is not shown on display.  Status indicates new image is being taken every 30 seconds and there are 36 images that haven't been downloaded.  Program would not respond to stop command.  Used Task Manager to stop TSI Manager.  Restarted program and raw and process images are now shown.  RTS at 1446 GMT.

9/4/08

PAD

Raw and processed images are not shown on monitor.  Task bar on TSI Manager shows:  Night  154.  Used Task Manager to stop TSI Manager program.  After restart, images were still not shown.  Stopped program and rebooted computer.  After restart, images are now shown and program is running correctly.  RTS at 1508 GMT.

8/6/08

PAD

TSI manager program is locked up.  Bottom of status bar shows:  Night  537.  No raw or processed images are shown.  Used Task manager to end program.  After restarting TSI manager, program started showing 537 images from 8/5/08 and then displayed current image.  RTS at 1542 GMT.

6/19/08

PAD

Mentor is on site today.  Shadowband is slightly out of alignment.  Mentor adjusted shadowband 4 deg. and slightly adjusted image alignment at 2033 GMT.

5/15/08

PAD

No images are shown on TSI Manager.  On status bar of window it shows Night and 709.  Used task manager to stop program and then restarted TSI Manager.  Program started and began to show the 709 images I assume were taken yesterday.  After that, current images were shown correctly.  RTS at 1536 GMT.

4/16/08

PAD

Computer is running but TSI program is not.  Program is not starting on its own in the mornings since computer was updated with security patches.  Notified Windows administrator who troubleshot issue with mentor.  At 1608 GMT, reset camera IP address.  RTS at 1455 GMT.

4/11/08

PAD

Display shows Windows Security Alert for file transfer.  TSI program is not running.  Notified Windows administrator who remotely accessed computer and checked settings for file transfer on firewall.  RTS at 1358 GMT.

3/11/08

PAD

Shadowband is shading but is close to needing alignment.  At 1524 GMT, adjusted shadowband 3 degrees and checked shading.  RTS at 1549 GMT.

2/27/08

PAD

Mirror has some chips missing on chrome surface.  At 1633 GMT, R/R mirror.  Reset 0 on mirror controller and checked shadowband alignment.  RTS at 1648 GMT.

2/4/08

PAD

Last SDS collection was on 2/1/08 at 2002 GMT.  Nothing is shown on TSI Manager window.  Scheduled tasks window is open and ftp_xfr file is highlighted. Printed screen of error and saved on desktop.  Stopped TSI Manager program, rebooted computer, and restarted program.  Program restarted correctly.  RTS at 1610 GMT.

12/24/07

MK

Due to a power outage the UPS has completely drained the internal batteries and is unable to restart. Reset UPS at 1430 GMT, all instruments in the Optical facility have been restarted and are currently running.

11/29/07

PAD

Last SDS collection was at 0402 GMT.  Found computer active and TSI manager display is up but nothing is shown on screen.  Computer is not responsive and program will not close.  Used Task Manager to end program.  Restarted program and display started showing images captured on 11/28/07.   After 11/28/07 images were displayed, current image was shown.  RTS at 1540 GMT.

9/21/07

PAD

Camera IP address needs to be reset.  Reset camera IP address as instructed by mentor.  Restarted TSI manager.  RTS at 1430 GMT.

8/30/07

PAD

Support bracket for camera arm was not sent with instrument after being repaired.  Bracket came in and needs to be installed.  Stopped image collection at 1752 GMT.  Installed bracket on camera support arm.  Checked and aligned image.  RTS at 1817 GMT.

7/31/07

MK

The instrument needs to be sent off for factory repairs. Removed instrument at 1500 GMT.

7/19/07

PAD

Current images are not shown on TSI Manager screen.  Task bar shows 319 images.  Program would not close.  Used Task Manager to end program.   Rebooted computer and restarted TSI Manager.  Program restarted correctly.  RTS at 1453 GMT.

7/18/07

PAD

Shadowband on mirror is shading but needs adjustment.  Adjusted shadowband 2 degrees and reset 0.  RTS at 1706 GMT.

6/11/07

PAD

Error displayed:  Image Capture in Progress.  Image capture will be terminated.  Cancel or Ok.   TSI manager will not bring up alignment image to check alignment.  Taskbar shows 4958 images.  TSI manager program will not close.  Used Task Mgr to end TSI manager program.    Rebooted computer and restarted TSI manager program.  Program started filing 4958 images.  Can now check alignment.  RTS at 1437 GMT.

6/4/07

PAD

TSI program locked sometime on 6/1/07.  At 1517 GMT today, observed TSI manager window is up but no images are displayed.  Used Window task manager to stop program.  Restarted program seems to be running correctly.  RTS at 1532 GMT.

5/29/07

MK

The power at the CF had failed some time over the weekend causing the UPS to shutdown due to low voltage on batteries. powered on UPS at 1434 GMT, restarted TSI computer and software at 1440 GMT.

5/14/07

PAD

Shadow band on mirror is shading sun but needs adjustment.    Could not connect computer with instrument using Hyperterminal.  Rebooted computer and was able to establish link.  Adjusted shadowband 3 deg. and reset 0.  RTS at 1440 GMT.

5/11/07

PAD

Error displayed:  Image Capture Error.  Printed screen of error message and saved in file on desktop.  Program will not restart.  Further troubleshooting by Tim Grove and Mark Klassen indicated computer is communicating with hub but not camera.  Notified instrument mentor who later returned call.  At 1720 GMT, troubleshot comm problem with mentor via phone.  Mentor reset IP address for camera.  Mentor also reset values displayed for temperature, lat & lon, etc. using mirtime HT.exe.  After IP for camera was reset, program restarted correctly.  RTS at 1759 GMT.

4/25/07

PAD

It is cloudy and unable to see if strip is shading correctly.  Alignment check shows mirror shading strip needs to be adjusted.  At 1535 GMT, adjusted mirror -3 deg and reset 0.  Unable to check correct shading due to clouds but alignment check shows correct alignment.  RTS at 1550 GMT.

4/18/07

PAD

Sun strip is shading but needs adjustment.  At 1440 GMT, adjusted sun strip on mirror 3 deg and reset 0.  Sun strip is not in correct alignment.  RTS at 1455 GMT.

4/11/07

PAD

TSI shadowband strip is still shading but needs adjustment.  At 1538 GMT, adjusted shadowband -3 deg. and reset 0.  Checked alignment.  RTS at 1553 GMT.

3/21/07

PAD

Shadowband is shading but needs adjustment.  Adjusted shadowband at 1538 GMT.  Verified shadowband is correctly aligned.  RTS at 1553 GMT.

3/20/07

PAD

Program is shown but is not running.  Status shows image capture error.  CF has experienced power bumps this morning.  Stopped and restarted TSI manager program.  Same error message is displayed.  Power cycled instrument and restarted program.  Program restarted correctly.   RTS at 2019 GMT.

3/20/07

PAD

Storms at CF tripped UPS located in Optical trailer.  Last successful collection was at 0300 GMT.  Following instruments were affected:  AERI01, MWR, RSS, SWS, TSI-880, SORTI01, and VCEIL25K.  Facilities reset UPS at 1600 GMT.  Restarted computers that did not restart after power was restored.  RTS at 1635 GMT.

3/14/07

PAD

At 1428 GMT, TSI manager only shows partial screen.  Instrument status shows Night.  Computer clock is correct.  TSI Manager window will not close.  Opened task manager and tried to close TSI manager.  Process would not end.  Chose end now option and process stopped.  Rebooted computer and restarted program.  RTS at 1443 GMT.

2/26/07

PAD

Shadowband is not shading sun.  Connected computer with instrument using Hyperlink.  Adjusted shadowband 7 deg. and reset 0, (home position).  Moved shadowband into correct position to check alignment.  Shadowband is aligned and shading sun.  RTS at 1632 GMT.

2/5/07

PAD

During PM rounds, found threaded collar that attaches mirror to drive motor had broken loose from mirror.  Mirror was still sitting on drive and was correctly shading.  R/R mirror with a used mirror in stock at the CF.  Aligned shadowband on mirror to 0 deg and reset.  Checked alignment of shadowband with sun.  Alignment checks OK.  Notified instrument mentor.  RTS at 1645 GMT.

6/14/06

PAD

Birds are landing on instrument enclosure and getting on mirror.  Mentor requests spike strips installed on corners of instrument enclosure surrounding mirror.  Installed spike strips on enclosure as instructed at 1510 GMT.  Notified instrument mentor.  RTS at 1525 GMT.

5/22/06

PAD

At 1537 GMT, observed TSI manager was running but no images were displayed.  Rebooted computer and restarted TSI manager.  Opened files and found last image captured was on 5/21/06 at 0129:30 GMT.  Program restarted correctly.  RTS at 1552 GMT.

5/16/06

PAD

Instrument mentor requests bird repellent spike strips installed on camera housing and arm.  At 1430 GMT, installed spike strips as requested.  RTS at 1445 GMT.

4/13/06

PAD

Instrument mentor filed a DQPR for instrument stating sky image is slightly out of focus.  Requested camera lens adjustment.  Chris Martin and I stopped image collection program at 1909 GMT.  Adjusted camera lens as instructed by mentor.  RTS at 1939 GMT. 

3/6/06

PAD

Raw and processed images are not displayed on computer.  Printed screen and saved image in my documents on TSI computer.  Rebooted computer at 1522 GMT.  Images are now displayed correctly.  RTS at 1537 GMT.

1/10/06

PAD

CF received 6" of snow last night.  Snow melted off mirror then froze mirror in place.  Mirror was unable to turn to align shading band.  At 1649 GMT, removed ice from mirror.  Mirror moved to correct shading alignment.  RTS at 1704 GMT.   

11/15/05

PAD

Display has TSI manager window open, no image shown.  Computer clock is correct.  Restarted computer at 1554 GMT.  Program started correctly.  RTS at 1609 GMT.

11/1/05

PAD

At 1606 GMT, raw image display is solid green and shows 100% cloud cover.  Stopped program and checked instrument.  Could not tell if shadowband is directly aligned with sun.  Opened instrument box and light in main power switch was not lit.  Power supply light is on.  Power cycled instrument and restarted program.  Instrument main power switch is still not lit.  Image is now displayed and appears normal.  RTS at 1621 GMT.

10/12/05

PAD

Instrument mentor is on site.  Mentor stopped image collections at 1429 GMT and adjusted shadowband –1 degrees and image alignment.  Started image collections at 1437 GMT.  RTS at 1444 GMT.

8/3/05

MR

The instrument mentor want to adjust the sky filter again. Adjusted the sky filter through the software control.  RTS at 2214 GMT.

8/2/05

MR

The instrument mentor wants the shadow band alignment adjusted through the software. Adjusted the band alignment with the software.  RTS at 1855 GMT.

8/2/05

MR

The instrument mentor wants to do a sky filter adjustment. Adjusted the sky filter through the software.  The filter will need to be readjusted again during the week.  RTS at 1840.

8/1/05

MR

The instrument mentor wants to do a band alignment and software adjustment. Performed the band alignment procedure.  Aligned the image through the software and focused the camera.  Started image collection at 2135 GMT.

7/28/05

MR

The TSI has arrived on site and needs to be reinstalled.  Reinstalled the TSI.  Could not get the camera to talk to the computer.  Tried to contact the mentor with no luck.  Left a message with the mentor and will await further instructions.  RTS at 1530 GMT.

6/7/05

PAD

Instrument mentor request instrument removed from service and shipped to him. Shut down computer and removed *TSI, (model 4404, *SN 100), at 1839 GMT.  Delivered instrument to shipping department.  RTS at 1855 GMT.

4/18/05

PAD

Last successful collection was on 4/16/05 at 0402 GMT.  Error displayed:  Hyper terminal applet has encountered a problem and needs to close.  Cleared error and restarted program.  Program restarted correctly.  Notified instrument mentor.  RTS at 1410 GMT.

4/14/05

V. Morris

Instrument mentor on site for maintenance activities.  Adjusted sky filter values from 1525 to 1535 GMT.  Leveled camera and adjusted center of image from 1735 to 1752 GMT.

4/15/05

PAD

Last successful collection was at 0400 GMT.  Error displayed:  Hyper terminal applet has encountered a problem and needs to close.  Cleared error and restarted program.  Program restarted correctly.  RTS at1441 GMT.

4/1/05

MR

The TSI 440 has arrived on site and needs to be installed.  Installed the instrument.  Due to the instrument not being able to communicate over the network, Tim grove will have to help troubleshoot when he returns on 4/4/05.  RTS at 2100 GMT.

3/22/05

MR

The instrument mentor requested that the TSI be checked due to communications with the instrument has been lost.  T/S found that the power supply monitor wire has come loose and has caused the software to think there is no power to the instrument.  Reconnected the monitor wire and communications was restored.  Notified the mentor.  The mentor thinks that the TSI will have to be replaced with a spare unit due to the LINIX program seems to have been corrupted.  Will await further instruction from the mentor.  This instrument will be down until a replacement unit has arrived.

3/21/05

MR

The TSI-880 power supply has failed and needs to be replaced.  Replaced the TSI power supply.  The TSI would power up but would not rotate to the proper position.  Contacted mentor and he said he would work on it remotely.

3/18/05

MR

The instrument will not respond by pinging or access by the web.  T/S found that the power lights for the imager, computer, 12 volt and 5 volt power supplies are not lit.  The power supply fuses that could be checked checked good but there is another fuse inside the power supply that could be bad.  The whole instrument has to be disassembled to get to this fuse.  Due to time constraints today, the instrument will be powered down and further T/S will be done on Monday.

2/28/05

PAD

Instrument mentor request instrument power cycled once/ month to reboot CPU. Shut down instrument at 1543 GMT.   Powered up instrument at 1603 GMT.  RTS at 1603 GMT.

1/31/05

PAD

Instrument mentor requests instrument power cycled once/ month to reboot CPU. Shut down instrument at 1540 GMT.  Restarted instrument at 1546 GMT.  RTS at 1555 GMT.

12/23/04

PAD

Instrument mentor requests instrument power cycled once/ month to reboot CPU. Shut down instrument at 1604 GMT.  Restarted at 1609 GMT.  RTS at 1619 GMT.

11/29/04

PAD

Instrument mentor requested instrument power cycled once/ month to reboot CPU.  Shut down instrument at 1613 GMT.  Restarted instrument at 1618 GMT.  RTS at 1628 GMT.

10/29/04

PAD

Instrument mentor requests instrument power cycled once/ month to reboot CPU. Shut down instrument at 1427 GMT.  Restored power at 1433 GMT.  Made sure instrument restarted correctly.  RTS at 1442 GMT.

10/1/04

PAD

Instrument mentor requested instrument power cycled once/ month to reboot CPU.  Shut down instrument at 1448 GMT.  Restored power at 1455 GMT.  Made sure instrument restarted correctly.  RTS at 1503 GMT.

8/30/04

PAD

Instrument mentor requested instrument power cycled once/ month to reboot CPU.  Shut off instrument at 1527 GMT.  Restored power at 1532 GMT.  Made sure instrument restarted correctly.  RTS at 1542 GMT.

6/30/04

MR

The power to the site was lost at 1330 GMT.  While power was out, the system was physically turned off so a controlled power up could be performed.  Power was restored to the site at 1455.  Powered up the system at 1545 GMT.  System came up and rebooted ok.  RTS at 1600 GMT.

6/29/04

PAD

Instrument mentor requested instrument power cycled once/ month to reboot CPU.  Shut down instrument at 1513 GMT.  Restored power at 1518 GMT.  Made sure instrument restarted correctly.  RTS at 1528 GMT.

6/2/04

PAD

Instrument mentor requested instrument power cycled once/ month to reboot CPU.  Shut off power to instrument at 1432 GMT.  Restored power at 1438 GMT.  Made sure instrument restarted correctly.  RTS at 1447 GMT.

4/29/04

PAD

Instrument mentor requested instrument power cycled once/ month to reboot CPU.  Shut down instrument at 2027 GMT.  Restored power at 2032 GMT.  Made sure instrument restarted correctly.  RTS at 2042 GMT.

4/1/04

PAD

Instrument mentor has requested instrument power cycled once/ month to reboot CPU.  Shut off instrument at 1545 GMT.  Powered up instrument at 1559 GMT.  Made sure instrument started correctly and display was shown.

2/23/04

PAD

Instrument mentor requested instrument be power cycled once/ month to reboot CPU.  Shut power off at 1610 GMT.  Started instrument at 1615 GMT.  Made sure instrument powered back up correctly and band shaded sun.  RTS at 1625 GMT.

1/26/04

PAD

Instrument mentor requested TSI-880 be power cycled once/ month to reboot CPU.  Shut off power to instrument at 1922 GMT.  Powered up instrument at 1928 GMT.  RTS at 1937 GMT.

12/29/03

PAD

Instrument mentor has requested TSI be power-cycled once/month to reboot CPU.  Shut off TSI instrument at 1656 GMT.  Powered up TSI at 1701 GMT.  Made sure display came back normally and band shaded sun.  RTS at 1711 GMT.        

11/24/03

PAD

Instrument mentor has requested TSI instrument be power cycled on monthly basis to reboot the CPU.  Power off the TSI instrument at 1522 GMT.  Restored power to instrument at 1530 GMT.  Checked display to make sure instrument restarted normally.  RTS at 1537 GMT.

10/27/03

PAD

Request from instrument mentor to power cycle the TSI to reboot the CPU.  Power cycled the TSI at 1640 GMT.  Power up TSI at 1645 GMT.  Made sure program was running.  RTS at 1655 GMT.

7/30/03

MR

I noticed that the heater for the mirror was off.  Turned on the mirror heater.  This will help dry out the mirror after a rain event.  RTS at 1500 GMT.

6/30/03

Mentor

The TSI shadow band was not completely centered.  Adjustment requires a username and password.  Mentor logged in to the TSI remotely and adjusted the shadow band by 6 degrees at 2347 GMT. 

6/25/03

MR

T/S found that the TSI had a bad connection inside the power supply for the TSI.  R/R the bad solder joints and tested the instrument.  The instrument seemed to work good.  Replaced the instrument back in the field and notified the mentor.  RTS at 1915 GMT.

6/17/03

JEM

Further troubleshooting of the TSI has found a power supply failure.  The TSI needs to be sent in for repairs.  Removed the TSI @1545 GMT to be prepared for shipment to be repaired. 

6/13/03

MR

The TSI is not responding to the web or to the laptop computer.  Yankee suggested that a monitor be connected to the unit and see if the computer is able to go through the boot process.  Connected a monitor to the unit and monitored the boot process.  The instrument was able to boot properly and worked for approx. 3 min then shut itself down.  Suspect the power supply is shutting down.  Notified Yankee and will await their response.

6/12/03

MR

The TSI is not responding to the web or a laptop computer.  Connected the laptop computer up to the instrument and I was not able to get a response from the unit.  Notified Yankee and will await their response.

6/9/03

JEM

The mentor suggests instead of just powering off the TSI, power it off and disconnect the DC battery to be 100% sure the computer inside is shutting down.  Powered off the TSI and disconnected the DC battery @1854 GMT.  Powered the TSI on and connected the DC battery @1900 GMT.  Both the 5V Power and Imager Power lights are on now.  Was able to connect to the TSI via web browser, but when the page was refreshed, the page could not be displayed.  Upon checking the TSI, the Imager Power light was no longer on.  Powered the TSI off and disconnected the DC battery again @1904 GMT.  Powered the TSI on and connected the DC battery @1910 GMT.  Both lights are lit again and the DSM activity light is going on and off.  Connected via web browser.  Refreshed the page several times with no problem.  RTS @1920 GMT.

6/9/03

JEM

The SDS has not collected from the TSI since Sat, Jun 07, 2003 0102 GMT.   Also am not able to connect to the TSI via a web browser.  Powered off the TSI for approximately 20 minutes.  Powered the TSI back on @1534 GMT.  Attempted to connect to the TSI with a web browser.  The attempt was unsuccessful.  The 5V power light was observed on, but the Imager Power light is not lit.  Checked all power and data connections inside the TSI.  None were found loose.  Powered off the TSI again @1540 GMT.  Powered on the TSI @1605 GMT.  Tried to connect via web browser @1610 GMT, but the attempt was again unsuccessful.  Will contact the mentor about this issue.  No further action at this time.  Did not RTS.

6/6/03

MR

The computer was locked up.  Power cycled the instrument so the computer would reboot.  The computer rebooted status OK.  RTS at 1453.

5/17/03

CAM

Remote monitoring of the DS View instrument status page indicated all of the instruments in the Optical trailer went down between 0300 and 0400 GMT on Saturday, 5/17.  T/S found that rodents had chewed through the MWR Profiler's power cables, which caused a direct short and killed the Optical trailer UPS.  This caused all the Optical trailer instruments to shut down.  Corrected the power short and brought the TSI back up at 1753 GMT.  RTS status O.K.

2/21/03

Mentor

Instrument mentor on site for maintenance.  Opened TSI and changed network settings at 2048 GMT.  Closed TSI and RTS at 2058 GMT.

2/20/03

Mentor

Mentor on site for maintenance/training.  Opened TSI for hardware tutorial and PM training at 1848 GMT.  Closed TSI and RTS at 1920 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

12/12/02

JEM

The TSI computer locked up on 12/12/02 @1334 GMT.  The computer was totally unresponsive to any keyboard commands.  Powered the computer off with the power switch @1444 GMT.  Powered the computer on @1445 GMT.  At 1553 the TSI computer was back online and the TSI was operating normally.  RTS @1448 GMT.

6/14/02

MR

Power restored.  System operational.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.  The instrument is shut down at this time.

6/5/02

CT

The camera lens was dirty and it needed to be cleaned. At 1437 GMT the TSI camera lens was cleaned. Status OK. RTS at 1452 GMT.

2/20/02

MR

The PC module and Power supply have been returned and need to be installed.  Installed both modules as per vendor’s instructions.  Powered up the instrument and observed the mirror did not rotate and would not track the sun as expected.  Notified the mentor and he notified Yankee.  Yankee contacted me and had me check the voltage in and out of the mirror control circuit.  The voltage in was observed to be 13.01 volts and there was no voltage out.  Yankee then had me connect a laptop to the instrument and observe the tracking algorithm but the algorithm was not present.  The mirror control circuit was suspected to have lost its programming during shipment.  Yankee will send a new motor control circuit over night express for us to install.

1/10/02

MR

The TSI is not able to be monitored over the network.  After numerous attempts to connect to the instrument via web or direct link with the laptop have failed, the instrument vendor has decided to have the PC module and power supply removed and shipped back to them.  Removed the modules and packed them up for shipping.  No serial numbers where observed on the modules.  RTS at 1630.

1/9/02

MR

The instrument mentor and the vendor want a monitor connected to the instrument so that the error message could be read.  Connected a monitor to the instrument and rebooted the system.  Upon reboot the system gave the following message: tsi 880.sgp.arm.gov login: Fatal server error: Could not create lock file in /tmp/.tx0-lock Sent the results tot he mentor and the vendor.  Awaiting reply.

1/8/02

MR

The instrument mentor wants the system power cycled due to it locking out remote access.  Power cycled the instrument at 1900 GMT.  This did not solve the remote access problem.  The instrument could not be locally accessed either.  Removed the PCMCIA storage card and power cycled again.  This did not cure the problem.  Connected a laptop computer to the instrument and tried to log on locally there but was unable to access the login screen after the password was entered.  Notified Chuck Long of the problem and will await his response.

1/8/02

CJ

I was unable to connect to the TSI on 1/8/2002. Power cycled instrument on 1/8/2002 at 1530 GMT. This did not correct the problem. Mentor has been informed of the problem. Trouble shooting is in progress. RTS 1/8/2002 at 1545 GMT.

12/17/2001

CJ

I was informed by Mike Rainwater that the TSI needed to be power cycled. Power cycled the TSI at 1557 GMT. At 1603 GMT power was restored to instrument. Status OK. RTS at 1612 GMT.

12/12/01

MR

The instrument mentor wants the camera from the old unit (Sn#102)) swapped out with the camera on the new unit (Sn#106).  Shut the unit down at 1505 GMT.  Removed the *camera from unit Sn# 106 and replaced it with the camera from unit Sn# 102.  Powered up the unit at 1520 GMT.  Called and e-mailed the mentor about the change so that he can notify Yankee Environmental and they can log onto the unit to reset the camera.  The mentor will notify me when this is done so the focus can be checked and adjusted.  RTS at 1530 GMT. 

12/5/01

MR

The instrument mentor wants TSI-102 removed and sent back to Yankee for upgrades.  He has sent tsi-106 for a temporary replacement.  Shut down the *TSI at 2030 GMT.  Removed the unit from its stand.  Unpacked and assemble unit 106 and placed it on the instrument stand.  Powered up the unit.  Following the instrument documentation, reset the IP address, DNS and gateway address' accordingly.  Powered cycled the instrument at 2210 so the new settings could take effect.  RTS at 2230.  Will pack instrument 102 up and send it back to Yankee for upgrades on RMA # 337.  *Old – 102  *New 106

11/26/01

CJ

The program was found locked up at 1545 GMT. Rebooted computer at 1545 GMT. RTS at 1600 GMT.

10/5/01

MR

The TSI would not communicate with the computer.  Rebooted the computer at 1447 but this did not fix the problem.  Rebooted the TSI at 1454 but this did not fix the problem.  T/S found that the network adapter had lost power due to the UPS it was plugged into had lost power.  After powering up the UPS and the network adapter the TSI communications was restored.  RTS at 1502 GMT.

9/17/01

MR

The instrument seemed to be working but could not been seen on the network.  Power cycled the instrument at 1425 GMT.  System came back up and was bale to be seen on the network at 1429 GMT.  RTS at 1440 GMT.

8/29/01

MR

There is a large sunflower plant starting to be seen in the field of view and needs to be removed. Removed the plant.  RTS at 1550.

8/27/01

MR

The TSI-880 stopped responding on Friday 8/24/01. Powered up the instrument.  Upon trying to hook a laptop up to the instrument it was discovered that the instrument seemed to be working.  Checked the network connection and it checked OK.  Checked the instrument with the laptop and it seemed OK.  The internal fan seemed to be surging but otherwise the instrument was working OK.  No further PM at this time.

8/13/01

MR

The TSI-880 is not tracking the sun.  Yankee systems contacted me and wants to check the boot sequence while connected to it with a laptop computer. Connected the laptop computer and watched the boot sequence.  The system booted ok and seemed to work.  The mirror stopped rotating and started tracking the sun.  Removed the cover on the camera.  I observed that the instrument seemed to work. Unhooked the laptop computer.  I was able to connect to the instrument via the web.  Will watch the system and see if it fails again.  Notified the mentor and Yankee system of the situation.  RTS at 1730 GMT.

8/9/01

MR

Morning rounds found that the TSI will not stop rotating.  Shut the system down and notified the mentor of the problem. After numerous times of power cycling and checking the fuses we shut the TSI down and notified the mentor.  The mentor notified Yankee Systems and we are to await instructions as to what could be wrong before performing any other maintenance.  Covered the camera lens.

7/31/01

MR

The TSI will not track the sun and cannot be accessed through the network. I Talked to Derrick Emond from Yankee environmental about the problem.  He had me check the 5 and 12 volt power supplies.  The 5 volt supply is ok but the 12 volt supply is reading 0.0 volts.  He suggested that we check the 5 amp fuse on the 12 volt supply.  After removing the chassis out of the instrument it was discovered that the 5 amp fuse was blown.  To keep from having to remove the chassis to check the fuse in the future I installed an inline fuse holder in this circuit so the fuse can be checked from above the chassis.  R/R the fuse and tested. The instrument tested good.  Checked the time, date and all the network and location settings and they checked good.  Reinstalled the chassis and restarted the instrument.  The instrument restarted ok and checked good.  RTS at 1730 GMT.

3/8/01

MR

A current sky image could not be seen when accessed over the network. Rebooted the TSI at 1623 GMT. After reboot the display was able to be seen. RTS at 1636.

2/9/01

MR

The mirror on the TSI was frozen up and would not move. The display was not able to see the sky. Checked the heater and found it was shut off. Turned on the heater and allowed the unit to warm up for about 10 mins. After 10 mins we were able to remove the ice on the mirror. The display returned to normal and the mirror was freed up. RTS at 1600 GMT.

2/6/01

MR

Due to this instrument being in the field since 6/30/00, the mentor wants the time and date checked for preventive measures. Checked the time and date using the procedure in the manual. The time looked as if it was off by 2 secs. Reset the time at 1621 GMT. RTS at 1630 GMT.

12/29/00

KT

Instrument mentor requested that the Bird-B-Gone be installed to prevent birds from landing on the instrument. At 1602 GMT removed tie wraps that were used until Bird-B-Gone was shipped in. At 1606 GMT installed one Bird-B-Gone on the arm of the instrument camera. At 1608 GMT installed two more Bird-B-Gone’ s on the top of the camera housing. RTS at 1615 GMT.

12/6/00

MR

The mentor wants us to try and place some sort of bird spikes on top of the camera housing and the horizontal camera bracket to help discourage the bird droppings on the mirror. Using tie wraps, I wrapped the ties around the horizontal camera bracket and left the wraps long and sticking up. Observed that the camera did not see them. RTS at 1630 GMT.

11/8/00

MR

Morning rounds found that the mirror on the instrument was icing over. This unit has a heater on the mirror that is supposed to prevent this. T/S found the heater was shut off. Turned the heater on but this overloaded the UPS that powers the TSI and the MWRC1. The UPS shut down and had to be restarted at 1638 GMT. Consulted with site facilities and found out that the TSI and the MWR heater power can be run on the same circuit. Shut the power down to the MWR and the TSI at 1830. Rerouted the power for the TSI from the UPS to the MWR heater circuit. This allowed the heater on the TSI to be powered on and not overload the UPS that powers the MWR. Powered up the circuit at 1840 GMT. Rebooted the TSI and the MWRC1.

10/30/00

MR

The mentor reports seeing fog in the data. The mentor wants us to check for moisture inside the electronics box. Checked for moisture and did not find any. Placed 2 large bags of desiccant inside to help absorb any unseen moisture.

10/19/00

MR

The TSI computer will not respond to the commands over the network. Power cycled the computer. The computer rebooted and everything checked O.K.

6/19/00

MR

The new disk drive came in on 6/14/00 and was installed that afternoon. The original problem of not being able to log into the unit over the network was solved but a new problem of no image was found. After numerous phone calls with the mentor it was determined that the imager network card is bad. A new card is being sent out today (6/19/00) and should be here on 6/20/00.

5/31/00

MR

Tim Grove and I where able to set up and connect the TSI to the network this morning. We where able to PING then instrument but not able to access it over the internet. Called the vendor and talked to Quentin Antes. After a few hours of working with it he called me back and said it looked as if the network hard drive has failed and wants it returned to him so it can be repaired or replaced. The RMA number for the return is 243. Shut the system down and removed the hard drive from the unit at 1930 GMT.

 5/25/2000

MR

Site Ops physically installed the TSI-880 and connected AC power. Will finalize installation and communications next week.